OpenStudy (anonymous):

Find the limit if it exists: (x^4-81)/(x-3) x->3

OpenStudy (anonymous):

\[x^4-3^4=(x^2+3^2)(x^2-3^2)=(x^2+3^2)(x-3)(x+3)\] So, what you really have is \[\frac{(x^2+3^2)(x-3)(x+3)}{x-3}=(x^2+3^2)(x+3) \;\; x\rightarrow3 \] \[(3^2+3^2)(3+3)=18*6=108\] and Bingo.

OpenStudy (anonymous):

hahahahhahahahhahahaha, how did you get yours to look like that?

OpenStudy (nottim):

I'm Batman.

OpenStudy (anonymous):

You can factor an (x-3) out of the numerator, which cancels out the (x-3) in the denominator. Then just plug in three to find the limit. This evaluates to 108.

OpenStudy (anonymous):

i thought you said earlier if the numerator has a higher power than the denom, then the limit doesn't exist?

OpenStudy (anonymous):

That only applies when x -> infinity

OpenStudy (anonymous):

Yes, that only applies when x->infinity. ^^ Tricks of the trade.